From Enrico, outlining tonights priorities:
The minimal set for the completion is to have similar measueremnt for the remaining points:
  • DONE15.8 <WFSmag < 16.5 Slow Bin 3
  • DONE16.51 < WFSmag < 17.0 Fast Bin 4
  • DONE17.0 < WFSmag < 18.0 Slow Bin 4
  • WFSmag > 18.0 (as far as we can go....)
For each point the procedure is the same used on the 24UT, with a reduced number of TNs.
- Preset with AOS in intervention mode and complete all the steps, including GOPT (enable NCPA for B3 only)
- Double check that the DIMM is properly measuring seeing conditions.
- Check the flux on LUCI (Ks), and find the minimal DIT providing a good signal (max 20s in order to avoid PSF drifts inthe DIT)
- Set the LUCI DIT in the Variable inspector of the AO system.
- Acquire the TN including the "PISCES PSF" with the required repetitions for a PSF that can provide a good measurement of the PSF halo.
- Collect at least 6 TNs for each star
- Offset out of the FoV to acquire SKY frames via the WFS control GUI with the same # of repetitions as in the TN

I remind here that performance measured with Seeing > 1.4" will not be useful for performance characterization.
So if seeing is steady above 1.4" above we can release the telescope for other ENG work.

I set the adsec limit to 26 deg for PEPSI obs at 23:22 but neglected to send out the email so Greg also set for good measure.

Pepsi observing until UT6:30 then handing over for SOUL Comm. Powered up Wavefront sensor in the afternoon.

5:45 noted that OCAM not live. Tried rebinning to no avail. Cycled fab 4: BCU interface, Slope Ctrl, OCAM2K controller, Master diagnostic and then rebinned and live again. Starting cooling.

6:45 RIP on center pupils. Ocam not live Corrumpted WFS. WFS Frame counter not incrementing.

071119 is junk, exposure time not set.

Slow bin4 Set

6:50 Star #12 , R= 17.1, WFS magnitude is 17.9. PA=0 rot 255
bin 4 OCAM frequency 139 Ocam gain 600 Average pupil intensity 15, 3.0 photons/subapp.

Optimize gain successful without safeskips

GOPT: optical gain 0.045 to start. variable GOPT because of low signal

Exptime 10x20s cube
DIMM 1.2" TN#20210303_071354 luci1..0088,
1.11" 20210303_071822 luci..0089,
0.97" 20210303_072232 luci1..0090,

Andrew offset us -20 20, Acquire Pisces dark: x10. TN#20210303_073343, luci1...0091, TN#20210303_073740, luci1...0092,

Andrew offset us back.

N/A 20210303_073927 luci1..0093,
N/A 20210303_074402 luci1..0094,

7:48 Star#19 PA0 R=16.0, no viable DIMM star in vicinity here.

8:02 Star #26. Dimm is struggling to acquire alpha boo....

R 16.0 WFSMag 17.2 Rotator 87. 0.087. This is not fast bin4. We are moving. Moon is likely contributing to background here.

Fast bin4 Set

8:10 Star 25 R=15.5. Rotator 98. WFS Estimated magnitude 16.9. Bin4 OCAM frequency 349 gain 600

tt2.01 ho10.79 ho20.10 Applying optical gain:0.091

180 in 5sec, taking 10x20sec cube again. we drifted off during that.

0.91" DIMM TN#20210303_082244 luci2.20210303.0097 Pupil error diverged pushing us off source. (GOPT died too as a result). camera lens X in limit. I had verified the rotator position during center pupils (as bin 3) but not during close loop as bin 4.

8:34 Resending with PA 60, rotator 161. WFS estimated magnitude 16.9

Apply optical gain 0.105 this time

Average pupil intensity 15, 2.9 photons/subapp.

10x20s cubes

DIMM0.92" TN20210303_084394 luci1..0099
DIMM1.05" TN20210303_084717 luci1..0100
DIMM1.02" TN20210303_085118 luci1..0101

Offset -20,20

2x20s test cube Pisces dark TN20210303_85647luci1..102

10x20s Pisces dark
DIMM1.06" TN20210303_090107 luci1..0103
DIMM1.05" TN20210303_090547 luci1..0104

offset back to source

DIMM 1.28" TN20210303_090629 luci1..0105
DIMM 1.02" TN20210303_091031 luci1..0106

Slow bin 3 Set

9:14 Star 24 R=14.9 WFS Estimate mag 16.0 Bin 3, frequency 352 gain 600. GOPT 0.111 NCPA enabled.
3.7 photons/subapp, 18 average pupil intensity, 174 rotator.

~2500count per 2.5sec exp. NEXP8sec. Taking 10x8sec cubes

DIMM 1.11" TN20210303_092510 luci1..0109
DIMM 1.23" TN20210303_092734 luci1..0110
DIMM 1.11" TN20210303_092938 luci1..0111

offset -20,20 for Pisces Dark - test dark 2x8s cube 1.09" TN20210303_93317 luci1..0112 ok 10x8s cube
DIMM 1.26" TN20210303_093535 luci1..0113
DIMM 1.23" TN20210303_093754 luci1..0114

Offset back

DIMM 1.01" TN20210303_093844 luci1..0115
DIMM 1.01" TN20210303_094600 luci1..0116
DIMM 1.01" TN20210303_094600 luci1..0117

Fast bin 3 Set

9:51 Star #23 R14.2. Estimated magnitude 15.4 on the wavefront sensor.. 3.1 photons/subapp. 15 average pupil intensity.
bin3 freq 760 gain 600 0.108. Optical gain 0.108. Taking test exp. Counts 3000 in teh peak in 2.5sec so taking 6 sec per dit

10x6s cube
DIMM 0.97" TN20210303_100138 luci1..0119
DIMM N/A" TN20210303_100445 luci1..0120
DIMM N/A" TN20210303_100635 luci1..0121

offset -20,20 for Pisces Dark - test dark 2x6s cube TN20210303_100941 luci1..0122 ok 10x6s cube
DIMM 1.10" TN20210303_101201 luci1..0123
DIMM 1.1" TN20210303_10... luci1..0124

offset back to source
DIMM 1.00" TN20210303_101434 luci1..0125
DIMM 1.1" TN20210303_101619 luci1..0126
DIMM N/A" TN20210303_101943 luci1..0127

Slow bin 4 Set

10:23 Sending prest to Star 33 R=16. Rotator at 158. Optical Gain Estimated magnitude 17.2 on the WFS. bin 4 OCAM freq 303
2.5 photons/subapp, 15 average pupil intesity

Enabled NCPA (habbit)

10x20s cube
DIMM 0.99" TN20210303_103228 luci1..0129.
DIMM 1.20" TN20210303_103657 luci1..0130 We walked off the patrol field

AOSGUI thinks adsec is off. Asdec will not open complaining that idl is still sending slopes. Restarting the AO arbitrator and adsec arb got things talking in the right way but not fixed.
Error while trying to save a tracking number froze up the WFS. Unable to open loop, tracking on sky for a bit there. Finally restarting the WFS and opening loop through teh freshly opened WFS resolved issue

relaod flat at 561 modes

Resending preset to Star 33 R 16 Estimated magntiude 17.2 bin4 OCAM Frquency 295, 0.087 optical gain

2.9 photons/subapp, 17 average pupil intensity

NO NCPA

DIMM N/A" TN20210303_110807 luci1..0132.
DIMM N/A" TN20210303_111229 luci1..0133.
DIMM 1.23" TN20210303_111706 luci1..0134.


offset -20,20 for Pisces Dark - test dark 2x20s cube TN20210303_112253 luci1..0135 ok
10x20s cube
DIMM 1.12" TN20210303_112739 luci1..0136. Offset placed the target in the field (-20,10)
Andrew took a test after moving the wrong way on the last offset and it looks like he left it with save mode normal....
DIMM 1.05" TN20210303_... luci1..0147 (offset -20,30) NORM SAVE MODE
DIMM 0.97" TN20210303_... luci1..0157 (offset -20,30) NORM SAVE MODE

Switching back to cube mode

Pisces dark try 1 million DIMM 1.24" TN20210303_114505... luci1..0159
DIMM 1.12" TN20210303_114905... luci1..0160

Back to source

DIMM 1.12" TN20210303_115006 luci1..0162
DIMM 0.91" TN20210303_115429 luci1..0163.

Slow bin 3 Set

Star 31 R=14.4 PA 60, rotator 183, Estimated Magnitude 16.1, bin 3,OCAM freq 347. Optical gain 0.112

Enable NCPA. 2.5 sec exp ~2100 so taking 10x8sec cube

DIMM 1.06" TN20210303_120819 luci1..0165.
DIMM 1.04" TN20210303_121034 luci1..0166.
DIMM 1.04" TN20210303_121238 luci1..0167.

Offset to sky (offset -20,20), taking test 2x8sec Pisces dark 01.06" TN20210303_121546 luci1..0168

Pisces dark 10x8s cube
DIMM 1.09" TN20210303_121838 luci1..0169. Safeskipping, opened
Reoptimizing and reapplying optical gains. NCPA reenabled
DIMM 1.01" TN20210303_122340 luci1..0170.
DIMM 1.05" TN20210303_122540 luci1..0171.

Offset back to source
DIMM N/A" TN20210303_122707 luci1..0172.
DIMM N/A" TN20210303_122944 luci1..0173.
DIMM 1.04" TN20210303_123150 luci1..0174.

Fast bin 3 Set

Star #22 R 13.2 PA 60 railed the camera lens in X at rotator position of 330 (aux loop gui Y). Sending a PA of 180 rotator 91. Estimated magntiude is 14.7 bin 3 916 Hz
Optical Gain 0.079

10x4s Cube

DIMM " TN20210303_125507 luci1..0176.
DIMM 1.22" TN20210303_125737 luci1..0177.
Offset to sky, Acquire Pisces Dark TN20210303_130132 luci1..0178
TN20210303_130248 luci1..0179

Warming OCAM, set Safe, power off. In idle. Done.

-- JenniferPower - 03 Mar 2021
Topic revision: r3 - 03 Mar 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback